4,11,11,6,13,7,4 Calculate the value of the sample variance. Round your answer to one decimal place.
Answer by Boreal(15235)

sum is 56
mean is 8
squared deviations are (absolute value used since squared)
4^2=16x2-32
3^2*2=18
2^2=4
5^2=25
1^1=1
sum=80
divide by 6=13.3 is variance